The current in self inductance L = 40 mH is to be increased uniformly from 1 amp to 11 amp in 4 milliseconds. The e.m.f. induced in the inductor during the process is


        (a)    100 volt    


(b)    0.4 volt


        (c)    4.0 volt


        (d)    440 volt


Grade:11

1 Answers

Hrishant Goswami
10 years ago

(a)

 

Given that L = 40 × 10-3 H,

di = 11 A – 1 A = 10 A and dt = 4 × 10-3 s